# PeakWeather: MeteoSwiss Weather Station Measurements for Spatiotemporal Deep Learning

### v1

@@ Acknowledgments
Appendix A Details about the NWP forecasting model ICON-CH1-EPS
The numerical weather prediction (NWP) baseline from the MeteoSwiss’ ICON-CH1-EPS consists of hourly forecasts at stations. For each forecast initialization every 3 hours, forecasts for 11 ensemble members are available. The first ensemble member is the control run, which has been initialized and run with unperturbed boundary conditions (i.e., the best estimate of the state of the atmosphere). The remaining 10 ensemble members use randomly perturbed boundary conditions and stochastically perturbed physics. This also implies that there is no connection between realizations of successive initializations other than for the control member and perturbed ensemble members should be treated as statistically exchangeable.
The forecasts cover the time span from forecast initialization to 33 hours in the future. For the aggregated quantities precipitation, sunshine, and wind_gust, the analysis timestep (i.e., forecast lead time 0) is missing.
The ICON-CH1-EPS forecasts at stations have been extracted from the values at the nearest grid point of the regular rotated latitude-longitude grid used at MeteoSwiss for processing of high-resolution NWP data; the distance between each station and its corresponding grid point is less than 1 km. To account for the difference in model topography and altitude, temperature forecasts from ICON-CH1-EPS are altitude-corrected. As is practice at MeteoSwiss, we use a constant lapse-rate of $0.06K/m$ for altitude correction. Relative humidity is computed from temperature and dew-point temperature via the ratio of saturated vapour pressure of water derived using Eq. 10 from (bolton1980computation, ). For the computation of relative humidity, neither temperature nor dew-point temperature is altitude corrected.
Appendix B Visualization of topographic descriptors
Figure 4 shows the topographic features included in PeakWeather, which are computed at both 2 km and 10 km spatial scales. The Aspect (Figures 3(a) and 3(b)) is defined as the azimuth (in degrees from North) of the steepest downslope direction. The Standard Deviation (STD) (Figures 3(c) and 3(d)) is the standard deviation of the surrounding elevation. The Topographic Position Index (TPI) (Figures 3(e) and 3(f)) describes whether a point is above or below the average elevation of its surroundings. The Slope (Figures 3(g) and 3(h)) is defined as the magnitude of the gradient vector of elevation. The West-East and South-North Derivatives (Figures 3(i), 3(j), 4(a), and 4(b)) are the partial derivatives of elevation in the west-east and north-south directions, respectively. Further information and implementation details can be found at the following link: https://github.com/MeteoSwiss/topo-descriptors.
@@ Appendix B Visualization of topographic descriptors
(i) WE-derivative 10km
(j) WE-derivative 2km
(a) SN-derivative 10km
(b) SN-derivative 2km
@@
Appendix C Variables availability over time
Figure 5 displays the number of stations per month that provide valid measurements for each of the 8 meteorological variables. The plot spans the entire dataset, from January 2017 to March 2025.
Figure 5: Stations with valid measurements of the meteorological variables included in PeakWeather.
@@ Appendix D Additional experimental details
Model architecture
The RNN and STGNN models have a 2-layer MLP as input encoder and 2-layer temporal processing via gated recurrent units with a hidden dimension of 128. The model predictions at each lead time and station are Gaussian distributions learned via the reparametrization trick, with mean and covariance conditioned on the inputs, lead time and station. The exponential linear unit is used as activation function, and 32-dimensional node embeddings are added as inputs to the encoder and readout layer, unless specified otherwise. STGNNs implement 6 message passing layers based on diffusion convolution from (li2018diffusion, ). Model hyperparameters have been selected as those the yielded minimum MAE on the validation set.
Graph construction
@@
The graph used in the experiments is constructed based on pairwise geographical distances between stations. The Gaussian kernel with width equal to the standard deviation of the station distances is applied to compute station similarity scores. To enforce sparsity in the graph, for each node, edges are included only for stations with a similarity above a predefined threshold of 0.7, and if the stations are among the $k=8$ nearest neighbors.
Model training and evaluation
@@
Model predictions are made for a 24-hour horizon from an input window of 24 time steps (1 day). All weather variables, encodings of the hour of the day and the day of the year, the binary mask of available observations, and all provided topographic descriptors are considered as input features to the model. Models are trained on the first 7 years of data by optimizing the Energy Score (ES) (gneiting2007strictly, ) at every station and every lead time (up to 24 hours ahead), validated on three months and tested on the last year, appropriately masking missing observations. The Energy Score is here estimated via Monte Carlo
@@
with ${\bm{y}}_{t+h}^{i}$ the target vector (wind velocity) and $\{\hat{\bm{y}}^{i,(m)}_{t+h}\}_{m=1}^{M}$ samples from the model’s predicted distribution. During training, $M=16$ samples are considered, while $M=100$ for testing. Training is performed for 200 epochs using Adam optimizer with an initial learning rate of 0.001 decreased at epochs 40, 60 and 120 by a factor of 0.3. Each epoch consists of 300 batches of size 32. Early stopping is triggered after 40 epochs without improvement on the validation forecasting MAE. Models are evaluated by averaging forecasting performance over the 24h horizon and evaluating at specific lead times. To assess the quality of wind direction forecasts, wind speeds below $1m/s$ are discarded.
Hardware and Software
Experiments are run on a machine with Intel(R) Xeon(R) CPU, 192 GB RAM, and NVIDIA L4 GPUs. The system ran Debian GNU/Linux 11. The code to run the experiments is developed in Python (rossum2009python, ), by relying mainly on the following open-source libraries: PyTorch (paske2019pytorch, ), PyTorch Lightning (Falcon_PyTorch_Lightning_2019, ), PyTorch Geometric (PyG) (fey2019fast, ), Torch Spatiotemporal (Cini_Torch_Spatiotemporal_2022, ), and the developed PeakWeather library (PeakWeatherBaseCode, ) to interface with the PeakWeather dataset (PeakWeatherDataset, ).
@@
The source code with all the configuration files necessary to reproduce the results is open-sourced on GitHub. [^2] https://github.com/Graph-Machine-Learning-Group/peakweather-wind-forecasting Model training and evaluation do not require particular hardware. Given the above machine and experimental settings, single model training and evaluation should complete within 3 hours.
Appendix E Illustrative forecast samples
@@
Figure 7 displays wind speed forecasts for 24h derived from the eastward and northward wind components for some sampled forecasting windows in the test set at different station locations. The figure compares the RNN-noemb and STGNN model predictions against those from ICON-CH1-EPS. Forecast uncertainty is also visualized.
@@
Figure 7: Comparison between RNN-noemb (left column), STGNN (center) and ICON-CH1-EPS (right column) for a selection of windows from the test set extracting the wind speed from the predicted eastward and northward wind components. For RNN-noemb and STGNN models, 50 predictive samples were produced, whereas ICON-CH1-EPS uses its 11 ensemble members. The colored solid line represents the sample median, while the black line represents the target quantity, obtained from the actual station measurements. Shaded areas denote forecast uncertainty, with the lighter region spanning from the 10th to the 90th percentile of the predictions, and the darker area highlighting the 25th to 75th percentile range.

### v2

@@ Acknowledgments
Appendix A Visualization of topographic descriptors and station distribution
Figure 3(a) shows the distribution of meteorological stations and rain gauges across the Swiss terrain. Figure 3(b) instead displays an example of a heuristically defined graph based on the distance between locations. Figure 3(c) shows the station elevation across longitude and latitude. Figure 5 shows the topographic features included in PeakWeather, which are computed at both 2 km and 10 km spatial scales. The Aspect (Figures 4(a) and 4(b)) is defined as the azimuth (in degrees from north) of the steepest downslope direction. The Standard Deviation (STD) (Figures 4(c) and 4(d)) is the standard deviation of the surrounding elevation. The Topographic Position Index (TPI) (Figures 4(e) and 4(f)) describes whether a point is above or below the average elevation of its surroundings. The Slope (Figures 4(g) and 4(h)) is defined as the magnitude of the gradient vector of elevation. The West-East and South-North Derivatives (Figures 5(a), 5(b), 5(c), and 5(d)) are the partial derivatives of elevation in the west-east and north-south directions, respectively. Further information and implementation details are provided by MeteoSwiss: https://github.com/MeteoSwiss/topo-descriptors.
Figure 3: Visualizations of the placement of PeakWeather stations. Panel 3(a)) distribution of the stations across the Swiss territory; circles denote meteorological stations, while triangles denote rain gauges. Panel 3(b)) a graph obtained using the geographical distance to compute a similarity heuristic for the meteorological stations. Panel 3(c)) vertical coverage of the stations.
@@ Appendix A Visualization of topographic descriptors and station distribution
(a) WE-derivative 10km
(b) WE-derivative 2km
(c) SN-derivative 10km
(d) SN-derivative 2km
@@
Appendix B Details about the NWP forecasting model ICON-CH1-EPS
The numerical weather prediction (NWP) baseline from MeteoSwiss ICON-CH1-EPS consists of hourly forecasts at stations. For each forecast initialization, occurring every 3 hours, forecasts for 11 ensemble members are available. The first ensemble member is the control run, which is initialized and run with unperturbed boundary conditions (i.e., the best estimate of the state of the atmosphere). The remaining 10 ensemble members use randomly perturbed boundary conditions and stochastically perturbed physics. This also implies that there is no connection between realizations of successive initializations other than for the control member and perturbed ensemble members should be treated as statistically exchangeable.
The forecasts cover the time span from forecast initialization up to 33 hours in the future. For the aggregated quantities precipitation, sunshine, and wind_gust, the analysis time step (i.e., forecast lead time 0) is missing.
The ICON-CH1-EPS forecasts at stations have been extracted from the values at the nearest grid point of the regular rotated latitude-longitude grid used at MeteoSwiss for processing of high-resolution NWP data; the distance between each station and its corresponding grid point is less than 1 km. To account for the difference in model topography and altitude, temperature forecasts from ICON-CH1-EPS are altitude-corrected. A constant lapse-rate of $0.06K/m$ is used for altitude correction. Relative humidity is computed from temperature and dew-point temperature via the ratio of saturated vapour pressure of water derived using Equation 10 from [33]. For the computation of relative humidity, neither temperature nor dew-point temperature is altitude corrected.
Appendix C Variable availability over time
Figure 6 displays the number of stations per month that provide valid measurements for each of the 8 meteorological variables. The plot spans the entire dataset, from January 2017 to October 2025.
Figure 6: Number of stations with valid measurements of the meteorological variables included in PeakWeather.
@@ Appendix D Additional experimental details
This appendix provides additional details about the experimental setup.
D.1 Model architectures
The RNN models are composed of a 2-layer gated recurrent unit (GRU) with a hidden size of 128 for the temporal processing. The TCN models are composed of 4 temporal convolution layers using a kernel size of 3, dilation of 2, and 128 hidden units. The MP-TCN and MP-RNN models extend TCN and RNN, respectively, by adding 6 graph message passing layers based on diffusion convolution [29], using a kernel size of 2. MP-TCN-gsl and MP-TCN+ variants include a graph structure learning module to respectively replace and augment the input graph used in MP-TCN with one learned from data. MP-TCN-gsl and MP-TCN+ use 2 message passing layers over the learned graph, while MP-TCN+ also includes 4 message passing layers over the original, proximity-based graph. For these models, hidden representations from the temporal processing and the message passing blocks are added together before applying the decoder.
The static node-level learnable embeddings ${\bm{e}}^{i}\in\mathbb{R}^{D_{e}}$ discussed in Tables 3 and 4 are learned end-to-end alongside the rest of the model parameters. As in [34], we use them to condition the encoder and decoder, both implemented as MLPs with exponential linear unit as activation function, to account for station-specific dynamics. Their size is set to 32. All models provide probabilistic forecasts as a Gaussian distribution (bivariate for wind forecasting and univariate for temperature forecasting, see Appendix F), for each lead time $h$, node $i$, and time step $t$. The Gaussian parameters are learned via the reparametrization trick and conditioned on the model inputs ${\bm{x}}_{t-W:t}^{i}$, ${\bm{u}}_{t-W:t}^{i}$, and ${\bm{v}}^{i}$. For PM-st and PM-day, the output distributions are $\mathcal{N}(\hat{\bm{y}}_{t+h}^{i}[d],\sigma_{t-W:t}^{i})$, with $\sigma_{t-W:t}^{i}$ being the sampling standard deviation estimated on the time window ${\bm{y}}_{t-W:t}^{i}[d]$.
These models are intentionally designed to be relatively straightforward deep learning models for temporal and spatiotemporal processing. The architectures and learning procedure are kept consistent across models. The main hyperparameters, such as the number of layers, the message passing operator, the activation, and learning rates are selected on the validation set.
Spatiotemporal models from the literature GraphWaveNet [31], AGCRN [30], and DCRNN [29] closely follow their original implementations and hyperparameters. The learning protocol and the probabilistic readout are the same as the one used for the other models. Chronos-2, a pre-trained 120-milion-parameter foundation model from [32], is selected for its supports of multivariate and covariate-informed forecasting. In our experiments, Chronos-2 is applied independently to each station, using the corresponding multivariate time series and associated covariates, with a substantially longer context window of 60 days that lets the model adapt to the temporal patterns of the specific timeseries. No fine-tuning was performed. Candidate window lengths ranging from 24 hours to 2 months were evaluated on the validation set. Unlike previous models, which are sample-based, Chronos-2 produces quantiles as outputs, therefore we consider the median to assess the MAE and do not report the Energy Score.
D.2 Graph construction and learning
@@
The heuristically defined graph used in the experiments is constructed based on pairwise geographical distances between stations. The Gaussian kernel with width equal to the standard deviation of the station distances is applied to compute station similarity scores. To enforce sparsity in the graph, each node is connected to at most its top $k=8$ nearest neighbors with a similarity score exceeding 0.7.
The graph structure learning module and training used in MP-TCN-gsl and MP-TCN+ models follow [35] to maintain computational efficiency throughout training and inference. The graph is modeled by matrix $\phi\in\mathbb{R}^{N\times N}$ of free model parameters that defines a distribution of top-$k$ graphs, with $k=10$. The parameters of $\phi$ are trained via score-function gradient estimation alongside the other model parameters. GraphWaveNet and AGCRN implement different graph structure learning strategies.
D.3 Model training and evaluation
@@
Model predictions are made for a 24-hour horizon from an input window of 24 time steps (1 day). All weather variables, encodings of the hour of the day and the hour of the year, the binary mask of available observations, and all provided topographic descriptors are considered as input features to the model. Models are trained on data up to September 2024 by optimizing the Energy Score (ES) [27] at every station and every lead time (up to 24 hours ahead) and tested on the last year, appropriately masking missing observations. Validated is performed on approximately one fifth of data covering October 2023 to September 2024. The Energy Score is here estimated via Monte Carlo
@@
with ${\bm{y}}_{t+h}^{i}$ the target vector (such as wind velocity) and $\{\hat{\bm{y}}^{i,(m)}_{t+h}\}_{m=1}^{M}$ samples from the model’s predicted distribution. During training, $M=16$ samples are considered, while $M=100$ for testing. When training MP-TCN+ and MP-TCN-gsl, a new graph is sampled for each batch, while, at inference time, the model relies on the mode of the distribution. Training is performed for 200 epochs using Adam optimizer [36] with an initial learning rate of 0.001 decreased at epochs 40, 60 and 120 by a factor of 0.3. Each epoch consists of 300 batches of size 32. Early stopping is triggered after 40 epochs without improvement on the validation forecasting MAE. Models are evaluated by averaging forecasting performance over the 24h horizon and evaluating at specific lead times. To assess the quality of wind direction forecasts, samples where the corresponding wind speed are below 1m/s are discarded.
As ICON issues predictions every 3 hours, we consider two test sets of data from October 2024. One consists only of the windows for which NWP forecasts are available and the other contains all possible windows. Unless specified otherwise, the former is considered to compare with ICON. Table 5 shows that the metrics are consistent across the two test sets.
D.4 Hardware and Software
@@
Experiments are run on a machine with Intel(R) Xeon(R) CPU, 192 GB RAM, and NVIDIA L4 GPUs. The system runs Debian GNU/Linux 11. The code to run the experiments is developed in Python [37], by relying mainly on the following open-source libraries: PyTorch [38], PyTorch Lightning [39], PyTorch Geometric (PyG) [40], Torch Spatiotemporal [41], and the developed PeakWeather library [12] to interface with the PeakWeather dataset [11]. The source code with all the configuration files necessary to reproduce the results is open-sourced on GitHub. [^2] https://github.com/Graph-Machine-Learning-Group/peakweather-forecasting Model training and evaluation do not require particular hardware. Given the above machine and experimental settings, single model training and evaluation should complete within 3 hours.
Appendix E Extended results for wind forecasting
To complement the empirical results of Table 3 and Figure 2, we report here additional results from the recurrent architectures and additional models from the literature.
Table 4 studies the conditioning of model predictions on station-specific components and relational information. Results for the RNN, MP-RNN, and GraphWaveNet confirm those in Table 3. In Table 4, GraphWaveNet-nogsl refers to the same GraphWaveNet architecture excluding the module that learns a graph in addition to the input one.
Table 5 instead compares the MAE on the velocity, speed, and direction, and Energy Score aggregated over the entire forecasting horizon. RNN-noemb refers to the RNN without learnable embeddings. Results confirm the superior performance of graph-based models in this setting and show consistent results across the two test sets.
Table 4: Impact on prediction performance of accounting for station-specific features and graph structures. The last columns indicate the size $D_{e}$ of the learnable node embeddings used (Emb.), and whether the corresponding model relies on station-specific topographic descriptors ${\bm{v}}_{i}$ (Topo.), an input graph (In.G.), and a graph structure learned from the data (GSL). Results are aggregated over five runs with different random seeds and are reported as mean $\pm$ standard deviation. Symbol “-” indicates non-applicable entries. For all metrics, the lower the better.
Model MAE Vel. MAE Speed MAE Dir. En.Score Emb. Topo. In.G. GSL
[…]
Table 5: Performance comparison of deep learning models with ICON and persistence model baselines for wind forecasting. Results report average $\pm$ 1 standard deviation computed over five runs with different random seeds, except for ICON and Chronos-2: ICON provides a single set of ensemble forecasts, while Chronos-2 predicted quantiles are not affected by initialization. As the persistence models require no training, their variability arises solely from Monte Carlo sampling during metric evaluation. The top part of the table considers the NWP test set with forecasts issued every 3h to compare with ICON, while the bottom part the full held-out test data.
Model MAE Velocity MAE Speed MAE Direction Energy Score Input Graph Learn Graph
NWP Test Set
[…]
ICON 1.132 1.113 39.073 1.407 - -
[…]
Chronos-2 1.108 1.202 42.96 - - -
Full Test Set
[…]
Appendix F Temperature forecasting: an inductive setup
We present an additional experiment on an inductive learning task for temperature prediction at previously unseen locations. When loading the dataset, both meteorological stations and rain gauges were considered. We consider all the stations equipped with a temperature sensor and do not use any covariates. Of those stations, 20 meteorological stations and 20 rain gauges are removed from the training set and used only for testing. This setting limits the breadth of models that can be tested, as graph structure learning and learned station embeddings cannot straightforwardly transfer to new nodes. Static attributes related to the station location, such as elevation and topographic features can instead be used. Besides these variations, the same hyperparameters used for wind forecasting are also considered here.
Table 6 summarizes the results in terms of MAE and ES for the various trainable models, persistence models, the foundation model Chronos-2, and ICON. Results are reported separately for each type of station (all, meteorological only, and rain gauge only) and further distinguished between results obtained on stations included in the training set (seen stations) and results obtained on the held-out set (unseen stations). First we comment that the performance is sensitive to the specific choice of held-out stations. We can see it from the difference in metrics between seen and unseen stations of the models of the first block (PM-st, PM-day, ICON, and Chronos-2), which are not trained on the seen/training stations; such sensitivity was observed—although not shown here—across different sets of held-out stations. For trained models (the second block with TCN, MP-TCN, and DCRNN) the predictions degrade slightly, overall by 10%. However, we note that for the spatiotemporal models, the performance is still significantly better than that of ICON and Chronos-2. Figure 7 illustrates the performance across lead times. Results suggest that local effects make this inductive setting challenging, requiring carefully designed models and training procedures, but the performance of graph-based models remains competitive nonetheless.
Figure 7: Performance comparison of deep learning models with ICON, persistence model baselines and Chronos-2 for inductive temperature forecasting. Results are averaged over five runs with different random seeds, except for ICON and Chronos-2. Shaded areas indicate $\pm$3 standard deviations. From left to right, the four plots represent the results obtained on seen meteorological stations, unseen meteorological stations, seen rain gauges, and unseen rain gauges.
Table 6: Performance comparison of deep learning models with ICON and persistence model baselines for temperature forecasting. Results report average $\pm$ 1 standard deviation computed over five runs with different random seeds, except for ICON and Chronos-2. For all seeds, the division seen-unseen stations is the same. The first part of the table reports the MAE, while the second part the Energy Score.
All stations Meteo stations Rain gauges
Model Seen Unseen Seen Unseen Seen Unseen
Mean Absolute Error (MAE)
[…]
Chronos-2 1.626 1.692 1.616 1.679 1.700 1.707
ICON 1.481 1.453 1.416 1.371 1.945 1.536
[…]
Energy Score (ES)
[…]
Appendix G Illustrative forecast samples
@@
Figure 8 displays wind speed forecasts for 24h derived from the eastward and northward wind components for some sampled forecasting windows in the test set at different station locations. These forecasts come from the experiment outlined in Section 5. The figure compares MP-TCN+ against the physics-based ICON model. Forecast uncertainty is also visualized as shaded areas, representing the the 5th, 25th, 75th and 95th quantiles computed from the samples. Similarly, Figure 9 displays the results of MP-TCN and ICON on the inductive learning task for temperature forecasting (Appendix F). The figure shows predictions at meteorological stations and rain gauges that were not part of the training set, illustrating the model’s ability to generalize to previously unseen locations.
@@
Figure 8: Visual comparison between MP-TCN+ (left column) and ICON-CH1-EPS (right column) for the wind forecasting task on the test set aligned with the ICON model. Wind speed is obtained as $\sqrt{u^{2}+v^{2}}$ from the eastward $u$ and northward $v$ wind components. For MP-TCN+, 100 predictive samples were produced, whereas ICON-CH1-EPS uses its 11 ensemble members. Shaded areas denote forecast uncertainty, in particular the 5th, 25th, 75th and 95th quantiles, as well as the median member in a solid line. The observed wind speed is displayed with a solid black line.
Figure 9: Visual comparison between MP-TCN (left column) and ICON-CH1-EPS (right column) for the inductive learning task of temperature forecasting. Results are shown for a subset of meteorological stations and rain gauges that were held out from the training dataset; MP-TCN was therefore not trained on observations from these locations.